Factor this polynomial correctly: 2x^2-22x-24
Answer by Theo(13342)

2x^2 - 22x - 24 = 2 * (x^2 - 11x - 12) = 2 * (x-12) * (x+1)

you confirm by multiplying the factors out to see if you get back to the original equation.

(x-12) * (x+1) is equal to x^2 + x - 12x - 12 which is equal to x^2 - 11x - 12.

(x^2 - 11x - 12) * 2 = 2x^2 - 22x - 24.

looks like we did good.

here's a lesson on how to factor that might be helpful.

http://www.purplemath.com/modules/factquad.htm